I need an equation and solution for this word problem(please show your work): four times the sum of three consecutive integers is 180.
Answer by stanbon(48545)

four times the sum of three consecutive integers is 180.
------------------------------
EQUATION:
4(x+(x+1)+(x+2)) = 180
3x+3 = 90
3x = 87
-------------
x = 29
x+1 = 30
x+2 = 31
-------------
Cheers,
Stan H.